# The extension of Schwarzschild line element to include uniformly accelerated mass

### v3

@@ Abstract
Our present work will find applications in various branches of Astrophysics and especially in the studies of gravitational wave and its detection.
Keywords: Gravitation – Spacetime metric – Schwarzschild object – accelerated motion – Rindler transformation.
@@ 1 Introduction
In general relativity the gravitational field of a static or a stationary object and the dynamics of the particles in such a field have been studied extensively in the past (e.g. [1, 2, 3, 4] etc.). But not much work has been done where the line element due to a moving gravitational mass has been calculated without any PPN approximation. So, it remains as one of the fascinating and hard subjects in the field of theoretical analysis. Although the study of the gravitational field due to moving objects is quite complicated, in a few cases e.g. [5, 6, 7, 8, 9, 10] the authors adopted various procedures to investigate the gravitational field of moving objects and suggested the dynamics of the particles in those fields. One of the interesting methods that some authors have used is the method of coordinate transformation to obtain the metric for moving objects. Guan-Sheng and Wen-Bin [11] obtained the spacetime line element for a uniformly moving point mass by performing the Lorentz transformation. Using the same approach, the gravitational field for moving Reissner-Nordström (R-N) black hole and moving Kerr black hole has been studied in [12] and [13] respectively. However, according to our knowledge, no work has been done on the gravitational field due to uniformly accelerated or arbitrarily accelerated gravitating objects.
@@
We organize our paper as: In Section 2 we investigate the gravitational field due to a uniformly moving Schwarzschild object. Consequently, we obtained the line element for such an object in terms of spatial Cartesian coordinates. Similarly in Section 3 we calculated the spacetime line element due to a uniformly accelerated Schwarzschild object along both $X$-direction and also along some arbitrary direction on the $XY$ plane. At last, in Section 4 we discuss the results obtained in the above calculations and conclude our findings.
2 Metric for a uniformly moving Schwarzschild object
@@
where $m$ is the mass parameter which is equal to half of the Schwarzschild radius $r_{g}=2GM/c^{2}$, $c$ is the speed of light in vacuum, $G$ is the gravitational constant, and $M$ is the mass of the object. The coordinate $t$ represents the time, and the spatial coordinates $(r,\theta,\phi)$ coincide with the ordinary spherical polar coordinates. Now, we consider that the Schwarzschild object is moving with uniform velocity along the $X$-direction. Then for the sake of convenience in studying the gravitational field, we rewrite the line element (1) in terms of the usual primed coordinate frame $K^{\prime}$ used as a moving frame in special relativity. So we re-write equation (1) as:
$ds^{2}$ $=\left(1-\frac{2m}{r^{\prime}}\right)c^{2}dt^{\prime 2}-\left(1-\frac{2m}{r^{\prime}}\right)^{-1}dr^{\prime 2}-r^{2}d\theta^{\prime 2}-r^{\prime 2}\sin^{2}\theta^{\prime}\,d\phi^{\prime 2}$ (2)
The spherical co-ordinates used above can be written in terms of the Cartesian co-ordinates as below :
$r^{\prime}$ $=\sqrt{x^{\prime 2}+y^{\prime 2}+z^{\prime 2}},\hskip 19.91684pt\theta^{\prime}=\cos^{-1}\frac{z^{\prime}}{\sqrt{x^{\prime 2}+y^{\prime 2}+z^{\prime 2}}},\hskip 19.91684pt\phi^{\prime}=\tan^{-1}\frac{y^{\prime}}{x^{\prime}}$ (3)
@@
Consequently, using the relation (3) we obtain the Schwarzschild line element in the $K^{\prime}$ coordinate frame as follows:
$ds^{2}$ $=\left[1-\frac{2m}{\sqrt{x^{\prime 2}+y^{\prime 2}+z^{\prime 2}}}\right]c^{2}dt^{\prime 2}-\Bigg{[}\Bigg{\{}\left(x^{\prime 2}+y^{\prime 2}+z^{\prime 2}\right)^{3/2}-2m\left(y^{\prime 2}+z^{\prime 2}\right)\Bigg{\}}dx^{\prime 2}$
$-\Bigg{\{}\Big{(}x^{\prime 2}+y^{\prime 2}+z^{\prime 2}\Big{)}^{3/2}-2m\left(x^{\prime 2}+z^{\prime 2}\right)\Bigg{\}}dy^{\prime 2}-\Bigg{\{}\left(x^{\prime 2}+y^{\prime 2}+z^{\prime 2}\right)^{3/2}$
$-2m\left(x^{\prime 2}+y^{\prime 2}\right)\Bigg{\}}dz^{\prime 2}-4mx^{\prime}y^{\prime}dx^{\prime}\,dy^{\prime}-4my^{\prime}z^{\prime}dy^{\prime}\,dz^{\prime}-4mx^{\prime}z^{\prime}dz^{\prime}\,dx^{\prime}\Bigg{]}$
$\times\Big{\{}\left(x^{\prime 2}+y^{\prime 2}+z^{\prime 2}\right)\left(\sqrt{x^{\prime 2}+y^{\prime 2}+z^{\prime 2}}-2m\right)\Big{\}}^{-1}$ (4)
Then we consider that the coordinate frame $K^{\prime}(ct^{\prime},x^{\prime},y^{\prime},z^{\prime})$ is moving with respect to the observer’s rest frame $K(ct,x,y,z)$ with uniform velocity $v_{x}$ in $X$-direction. Additionally, we also consider that for the frames $K$ and $K^{\prime}$ the respective $X$ and $X^{\prime}$ axes are coinciding and other coordinate axes $(Y,Z)$ and $(Y^{\prime},Z^{\prime})$ are parallel to each other. Meanwhile, both the frames of reference $K$ and $K^{\prime}$ are inertial frames, and hence the relations between the coordinates of two such frames can be established using the Lorentz transformation. Therefore using the Lorentz transformation one can write [21]:
$ct$ $=\gamma\left(ct^{\prime}+\frac{v_{x}}{c}\cdot x^{\prime}\right),\hskip 28.45274ptx=\gamma\left(x^{\prime}+v_{x}\,t^{\prime}\right),\hskip 28.45274pty$ $=y^{\prime}\,\,,\hskip 28.45274ptz=z^{\prime}$ (5)
@@
$ct^{\prime}$ $=\gamma\left(ct-\frac{v_{x}}{c}\cdot x\right),\hskip 28.45274ptx^{\prime}=\gamma\left(x-v_{x}\,t\right),\hskip 28.45274pty^{\prime}$ $=y\,\,,\hskip 28.45274ptz^{\prime}=z$ (6)
@@
$cdt^{\prime}$ $=d(ct^{\prime})=\gamma\left(cdt-\frac{v_{x}}{c}dx\right)\text{\,\,\,\,or\,\,\,\,}c^{2}dt^{\prime 2}=\gamma^{2}\left[c^{2}(dt)^{2}+\frac{v_{x}^{2}}{c^{2}}(dx)^{2}-2\frac{v_{x}}{c}dxcdt\right]$ (7)
@@
Now following the similar approach used in [11, 12, 22] we obtain the line element by putting (6)–(9) in equation (4) and simplifying as below:
$ds^{2}$ $=\gamma^{2}\Bigg{[}1-\frac{2m}{\sqrt{\mathcal{R}}}+\frac{v_{x}^{2}}{c^{2}}\Bigg{\{}2m\left(y^{2}+z^{2}\right)-\Big{(}\mathcal{R}\Big{)}^{3/2}\Bigg{\}}\Bigg{\{}\Big{(}\mathcal{R}\Big{)}\Big{(}\sqrt{\mathcal{R}}-2m\Big{)}\Bigg{\}}^{-1}\Bigg{]}c^{2}dt^{2}$
$+\gamma^{2}\Bigg{[}\frac{2m\left(y^{2}+z^{2}\right)-\left(\mathcal{R}\right)^{3/2}}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}+\frac{v_{x}^{2}}{c^{2}}\left(1-\frac{2m}{\sqrt{\mathcal{R}}}\right)\Bigg{]}dx^{2}$
$+\Bigg{[}\frac{2m\left(\gamma^{2}(x-v_{x}t)^{2}+z^{2}\right)-\left(\mathcal{R}\right)^{3/2}}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}dy^{2}+\Bigg{[}\frac{2m\left(\gamma^{2}(x-v_{x}t)^{2}+y^{2}\right)-\left(\mathcal{R}\right)^{3/2}}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}dz^{2}$
$+\frac{4\gamma^{2}m\,v_{x}\Bigg{\{}\sqrt{\mathcal{R}}\left(\gamma^{2}(x-v_{x}t)^{2}+\mathcal{R}\right)-2m\left(\mathcal{R}\right)\Bigg{\}}}{\left(\mathcal{R}\right)^{3/2}\left(\sqrt{\mathcal{R}}-2m\right)}dx\,dt$
$+\Bigg{[}\frac{4\gamma^{2}my\,v_{x}(x-v_{x}t)}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}\,dydt+\Bigg{[}\frac{4\gamma^{2}mz\,v_{x}(x-v_{x}t)}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}dz\,dt$
$-\Bigg{[}\frac{4\gamma^{2}my(x-v_{x}t)}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}\,dydx-\Bigg{[}\frac{4myz}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}dy\,dz$
$-\Bigg{[}\frac{4\gamma^{2}mz(x-v_{x}t)}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}dz\,dx$ (10)
where $\mathcal{R}=\gamma^{2}(x-v_{x}t)^{2}+y^{2}+z^{2}$.
The line element (10) represents the gravitational field of a uniformly moving Schwarzschild object. Below we analyze this line element at different limiting conditions:
@@
    $ds^{2}$ $=\left[1-\frac{2m}{\sqrt{x^{2}+y^{2}+z^{2}}}\right]c^{2}dt^{2}-\Bigg{[}\Bigg{\{}\left(x^{2}+y^{2}+z^{2}\right)^{3/2}-2m\left(y^{2}+z^{2}\right)\Bigg{\}}dx^{2}$
    $-\Bigg{\{}\Big{(}x^{2}+y^{2}+z^{2}\Big{)}^{3/2}-2m\left(x^{2}+z^{2}\right)\Bigg{\}}dy^{2}-\Bigg{\{}\left(x^{2}+y^{2}+z^{2}\right)^{3/2}$
    $-2m\left(x^{2}+y^{2}\right)\Bigg{\}}dz^{2}-4mxydx\,dy-4myzdy\,dz-4mxzdz\,dx\Bigg{]}$
    $\times\Big{\{}\left(x^{2}+y^{2}+z^{2}\right)\left(\sqrt{x^{2}+y^{2}+z^{2}}-2m\right)\Big{\}}^{-1}$ (11)
    which is exactly similar to the Schwarzschild line element in Cartesian coordinates as shown in equation (4) as $ct=ct^{\prime}$, $x=x^{\prime}$, $y=y^{\prime}$ and $z=z^{\prime}$ for $v_{x}=0$.
@@
    Case-III: Following a similar approach used in [5, 23] we calculate a spacetime metric from equation (10) in which the velocity of the object is close to the speed of light i.e. $v_{x}\rightarrow c$. Here using the relativistic mass $M_{r}$ and energy $E$ relation: $E=M_{r}c^{2}$, we write $E=\frac{Mc^{2}}{\sqrt{1-\frac{v_{x}^{2}}{c^{2}}}}$ and hence obtain $m=\frac{GM}{c^{2}}=\frac{GE}{c^{4}}\sqrt{1-\frac{v_{x}^{2}}{c^{2}}}$, where $M$ is the rest mass and $G$ is the gravitational constant. The obtained spacetime metric for $v_{x}\rightarrow c$ can be written as (see Appendix A eq. (A.13) ):
@@
    which is exactly same as the Aichelburg-Sexl metric derived in [5, 23]. The difference is only in the choice of the reference frame to represent the equation (13) where the original Aicherburg-Sexl metric is written in terms of the primed coordinate frame (see Eqn. (A.1)). Therefore, we can state that the line element (10) reduces to the Aicherburg-Sexl metric under the condition $v_{x}\rightarrow c$.
Here we admit that, we are fully aware of the difficulties in writting Lorentz transformation equations in the background of Schwarzschild geometry. Paithankar and Kolekar [24] had discussed about a Schwarzschild body, in the Minkowski-Rindler setting such that the black hole is at a large distance away from the Rindler observer. The spacetime geometry in the local neighborhood of the Rindler trajectory will still be dominantly flat but now also has small perturbations to the background flat metric due to the presence of the black hole.
@@
In the contrary, in our present paper we took an alternate approach. We at first write the Schwarzschild line element in Cartesian coordinate form. Then we consider the frame in which the Schwarzschild object is placed, is undergoing uniform motion. Such form of the line element enables us to use the standard Lorentz transformation without any approximated transformation relations. So, in our case we used the standard Lorentz transformation in Cartesian form to obtain the line element of a moving Schwarzschild object. Although strictly speaking the specific properties of the parallel transport of vectors in Minkowski spacetime is not allowed in Schwarzschild geometry [24], but we believe the approximate solution that we shall obtain by our method, will be at least comparable to the other set of approximated results in [11]. The error in our approximation can be further reduced by considering a far field solution ( isotropic form) of Schwarzschild line element, which we postpone for future. We can say that our treatment can be considered as an alternate approach to this problem.
Since in our case we have used the standard Lorentz transformation relation in Cartesian form of the Schwarzschild line element, we don’t expect the result obtained in our paper to match exactly with the results obtained by the authors in [11] through harmonic coordinate transformation.

### v4

@@ Abstract
Keywords: Gravitation; Spacetime metric; Schwarzschild object; accelerated motion; Rindler transformation.
@@ 1 Introduction
In general relativity the gravitational field of a static or a stationary object and the dynamics of the particles in such a field have been studied extensively in the past (e.g. [1, 2, 3, 4, 5] etc.). But not much work has been done where the line element due to a moving gravitational mass has been calculated without any PPN approximation. So, it remains as one of the fascinating and hard subjects in the field of theoretical analysis. Although the study of the gravitational field due to moving objects is quite complicated, in a few cases e.g. [6, 7, 8, 9, 10, 11] the authors adopted various procedures to investigate the gravitational field of moving objects and suggested the dynamics of the particles in those fields. One of the interesting methods that some authors have used is the method of coordinate transformation to obtain the metric for moving objects. Guan-Sheng and Wen-Bin [12] obtained the spacetime line element for a uniformly moving point mass by performing the Lorentz transformation. Using the same approach, the gravitational field for moving Reissner-Nordström (R-N) black hole and moving Kerr black hole has been studied in [13] and [14] respectively. On the other hand, Paithankar and Kolekar [15] discussed about a Schwarzschild body, in the Minkowski-Rindler setting such that the black hole is at a large distance away from the Rindler observer. They also mentioned that, in this situation the spacetime geometry in the local neighborhood of the Rindler trajectory will be dominantly flat with small perturbations to the background flat metric due to the presence of the black hole. Later, they investigated the trajectories of linear uniform acceleration in the background of Schwarzschild geometry and compared their results with that of the $C$-metric. The $C$-metric is an exact solution of the Einstein field equation, which represents a pair of black holes with equal mass accelerated away from each other due to the presence of conical singularities [16, 17]. Studies on light-like geodesics and time-like geodesics in $C$-metric spacetime can be found in [18] and [19], respectively. However, to our knowledge, no work has been done using the coordinate transformation approach to investigate the gravitational field due to uniformly accelerated or arbitrarily accelerated gravitating objects except a few, as like in the case of uniformly moving black holes [6]. Therefore, it is necessary to study the gravitational field due to an accelerated object using the coordinate transformation method.
@@
We organize our paper as follows. In Section 2 we discuss the isotropic form of the Schwarzschild line element. Then, we obtain a far-field approximation of the Schwarzschild line element in the Cartesian coordinates. In Section 3 we investigate the gravitational field due to a uniformly moving Schwarzschild object. Consequently, we obtained the line element for such an object in terms of spatial Cartesian coordinates. Similarly in Section 4 we calculated the spacetime line element due to a uniformly accelerated Schwarzschild object along $X$-direction and also along some arbitrary direction on the $XY$ plane. Finally, in Section 5 we discuss the results obtained in the above calculations and conclude our findings.
2 Schwarzschild line element under far-field approximation using isotropic cartesian coordinate system
@@
where $m$ is the mass parameter which is equal to half of the Schwarzschild radius $r_{g}=2GM/c^{2}$, $c$ is the speed of light in vacuum, $G$ is the gravitational constant, and $M$ is the mass of the object. The line element (1) can be written in isotropic form by assuming a radial function $\rho$ which is related to the radial coordinate $r$ of the spherical polar coordinate as follows [2, 3]:
$r$ $=\left(1+\frac{m}{2\rho}\right)^{2}\rho$ (2)
@@
So using the relation (2), we can obtain the isotropic form of the Schwarzschild line element as below [2, 3]:
$ds^{2}$ $=\left(\frac{1-\frac{m}{2\rho}}{1+\frac{m}{2\rho}}\right)^{2}c^{2}dt^{2}-\left(1+\frac{m}{2\rho}\right)^{4}\Big\{d\rho^{2}+\rho^{2}d\theta^{2}+\rho^{2}\sin^{2}\theta\,d\phi^{2}\Big\}$ (3)
where $\rho$, $\theta$, $\phi$ are the isotropic spherical coordinates. Now we introduce isotropic Cartesian coordinates $x$, $y$, $z$ and further we apply far-field approximation, i.e. $\rho>>m$. Therefore we can approximately write from equation (3) as follows [2]:
$ds^{2}=\left(1-\frac{2m}{\rho}\right)c^{2}dt^{2}-\left(1+\frac{2m}{\rho}\right)\Big\{dx^{2}+dy^{2}+dz^{2}\Big\}$ (4)
Again, from equation (2) expanding $r/\rho$ in the power of $m/\rho$ we obtain as:
$\frac{r}{\rho}$ $\approx 1+\frac{m}{\rho}+\frac{m^{2}}{4\rho^{2}}$ (5)
Since $\rho>>m$, so neglecting the terms higher than $m/\rho$, we can obtain as:
$\rho$ $\approx r-m$ (6)
Then substituting from equation (6) into equation (4) we can obtain a far-field approximated Schwarzschild line element as:
$ds^{2}=\left(1-\frac{2m}{r-m}\right)c^{2}dt^{2}-\left(1+\frac{2m}{r-m}\right)\Big\{dx^{2}+dy^{2}+dz^{2}\Big\}$ (7)
where $r=\sqrt{x^{2}+y^{2}+z^{2}}$.
We use the line element (7) to study the gravitational field due to a Schwarzschild object undergoing uniform motion and accelerated motion.
3 Metric for a uniformly moving Schwarzschild object
Now, we consider that the Schwarzschild object is moving with uniform velocity along the $X$-direction. Then for the sake of convenience in studying the gravitational field, we rewrite the line element (7) in terms of the usual primed coordinate frame $K^{\prime}$ used as a moving frame in special relativity as follows:
$ds^{2}=\left(1-\frac{2m}{r^{\prime}-m}\right)c^{2}dt^{\prime 2}-\left(1+\frac{2m}{r^{\prime}-m}\right)\Big\{dx^{\prime 2}+dy^{\prime 2}+dz^{\prime 2}\Big\}$ (8)
where $r^{\prime}=\sqrt{x^{\prime 2}+y^{\prime 2}+z^{\prime 2}}$. Then we consider that the coordinate frame $K^{\prime}(ct^{\prime},x^{\prime},y^{\prime},z^{\prime})$ is moving with respect to the observer’s rest frame $K(ct,x,y,z)$ with uniform velocity $v_{x}$ in $X$-direction. Additionally, we also consider that for the frames $K$ and $K^{\prime}$ the respective $X$ and $X^{\prime}$ axes are coinciding and other coordinate axes $(Y,Z)$ and $(Y^{\prime},Z^{\prime})$ are parallel to each other. In curved spacetime (viz Schwarzschild) the relation between two such frames cannot be directly established by using the standard Lorentz transformation relations. It is because Lorentz transformation uses specific properties of the Minkowski spacetime such as the parallel transport of vectors allowed by its very particular affine structure, which the Schwarzschild geometry does not allow. However, in our present work, we considered a far-field condition where the spacetime geometry is close to flat, that can be described using the line element in equation (8). So in this context, we can use the Lorentz transformation relations written in terms of Cartesian coordinates as follows:
$ct$ $=\gamma\left(ct^{\prime}+\frac{v_{x}}{c}\cdot x^{\prime}\right),\hskip 14.22636ptx=\gamma\left(x^{\prime}+v_{x}\,t^{\prime}\right),\hskip 14.22636pty$ $=y^{\prime}\,\,,\hskip 14.22636ptz=z^{\prime}$ (9)
@@ 3 Metric for a uniformly moving Schwarzschild object
$ct^{\prime}$ $=\gamma\left(ct-\frac{v_{x}}{c}\cdot x\right),\hskip 14.22636ptx^{\prime}=\gamma\left(x-v_{x}\,t\right),\hskip 14.22636pty^{\prime}$ $=y\,\,,\hskip 14.22636ptz^{\prime}=z$ (10)
@@
$cdt^{\prime}$ $=\gamma\left(cdt-\frac{v_{x}}{c}dx\right),\,\,\,c^{2}dt^{\prime 2}=\gamma^{2}\left[c^{2}(dt)^{2}+\frac{v_{x}^{2}}{c^{2}}(dx)^{2}-2\frac{v_{x}}{c}dxcdt\right]$ (11)
@@
Now following the similar approach used in [6, 7, 12, 13] we obtain the line element by putting (10)–(13) in equation (8) and simplifying as below:
$ds^{2}$ $=\Bigg[1-\frac{2m\gamma^{2}}{\mathcal{R}}\Big\{1+\frac{v_{x}^{2}}{c^{2}}\Big\}\Bigg]c^{2}dt^{2}-\Bigg[1+\frac{2m\gamma^{2}}{\mathcal{R}}\Big\{1+\frac{v_{x}^{2}}{c^{2}}\Big\}\Bigg]dx^{2}$
$-\Big[1+\frac{2m}{\mathcal{R}}\Big]\Big\{dy^{2}+dz^{2}\Big\}+\frac{8\gamma^{2}mv_{x}}{\mathcal{R}}\,dxdt$ (14)
where $\mathcal{R}=\sqrt{\gamma^{2}\left(x-v_{x}\,t\right)^{2}+y^{2}+z^{2}}-m$. The line element (14) represents the gravitational field of a uniformly moving Schwarzschild object. Below we analyze this line element at different limiting conditions:
@@
    $ds^{2}$ $=\left[1-\frac{2m}{r-m}\right]c^{2}dt^{2}-\left[1+\frac{2m}{r-m}\right]\Big\{dx^{2}+dy^{2}+dz^{2}\Big\}$ (15)
    which is exactly similar to the Schwarzschild line element in Cartesian coordinates as shown in equation (8) as $ct=ct^{\prime}$, $x=x^{\prime}$, $y=y^{\prime}$, $z=z^{\prime}$ and $r=r^{\prime}$ for $v_{x}=0$.
@@
    Case-III: Following a similar approach used in [6, 27] we calculate a spacetime metric from equation (14) in which the velocity of the object is close to the speed of light i.e. $v_{x}\rightarrow c$. Here using the relativistic mass $M_{r}$ and energy $E$ relation: $E=M_{r}c^{2}$, we write $E=\frac{Mc^{2}}{\sqrt{1-\frac{v_{x}^{2}}{c^{2}}}}$ and hence obtain $m=\frac{GM}{c^{2}}=\frac{GE}{c^{4}}\sqrt{1-\frac{v_{x}^{2}}{c^{2}}}$, where $M$ is the rest mass and $G$ is the gravitational constant. Then substituting this expression for $m$ into equation (14) we obtain as:
    $ds^{2}$ $=\Bigg[1-\frac{2GE}{c^{4}\mathcal{R}_{0}}\Big\{1+\frac{v_{x}^{2}}{c^{2}}\Big\}\Bigg]c^{2}dt^{2}-\Bigg[1+\frac{2GE}{c^{4}\mathcal{R}_{0}}\Big\{1+\frac{v_{x}^{2}}{c^{2}}\Big\}\Bigg]dx^{2}$
    $-\Big[1+\frac{2GE}{c^{4}\mathcal{R}_{0}}\Big\{1-\frac{v_{x}^{2}}{c^{2}}\Big\}\Big]\Big\{dy^{2}+dz^{2}\Big\}+\frac{8GEv_{x}}{c^{4}\mathcal{R}_{0}}dxdt$ (17)
    where $\mathcal{R}_{0}=\Big\{\left(x-v_{x}\,t\right)^{2}+\Big\{1-\frac{v_{x}^{2}}{c^{2}}\Big\}(y^{2}+z^{2})\Big\}^{1/2}-\frac{GE}{c^{4}}\Big\{1-\frac{v_{x}^{2}}{c^{2}}\Big\}$. In the limit of $v_{x}\rightarrow c$, the above line element (17) can be written as:
@@
    which is exactly same as the Aichelburg-Sexl metric [^1] In the original derivation, Aichelburg and Sexl [6] considered $c=G=1$ and $E=p$ as a constant. derived in [6, 27]. The difference is only in the choice of the reference frame to represent the equation (18) where the original Aicherburg-Sexl metric is written in terms of the primed coordinate frame. Therefore, we can state that the line element (14) reduces to the Aicherburg-Sexl metric under the condition $v_{x}\rightarrow c$.
@@
In the contrary, in our present paper we took an alternate approach. We at first write a far-field approximated Schwarzschild line element in Cartesian coordinates. Then we consider the frame in which the Schwarzschild object is placed, is undergoing uniform motion. Such form of the line element enables us to use the standard Lorentz transformation relations. So, in our case we used the standard Lorentz transformation in Cartesian form to obtain the line element of a moving Schwarzschild object. Although strictly speaking the specific properties of the parallel transport of vectors in Minkowski spacetime is not allowed in Schwarzschild geometry [15], but we use the far-field approximated line element (7), which allow us to perform the Lorentz transformation in the present context as shown in [7, 10, 29].
Since in our case we have used the standard Lorentz transformation relation in the Cartesian form of the far-field approximated Schwarzschild line element, we don’t expect the result obtained in our paper to match exactly with the results obtained by the authors in [12] through harmonic coordinate transformation.